OpenStudy (laronjaxon):

4x/9+x=10/9 simplify! please help! show work! adding and subtracting rationals! pleaseee!

OpenStudy (anonymous):

start by multiplying everything by 9 this will get rid of all the denominators the equation turns into 4x+x=10 5x=10 x=2

OpenStudy (laronjaxon):

can you show me like in a full worked out problem so ican try to do the rest of these please :o?

OpenStudy (anonymous):

(2/5)x + (4/5) = 9 how do you get rid of the denominators??

OpenStudy (anonymous):

you would multiply everything by 5 (2/5) times 5 = 2 (4/5) times 5 = 4 9 times 5 = 45 2x+4=45 subtract 4 from both sides then divide both sides by 2 answer: x=20.5

OpenStudy (anonymous):

you multiply everything by the LCD of the fractions LCD = lowest common denominator for simple problems where all the denominators are the same. the LCD is that denominator for more complex ones, the LCD is the smallest number that all the denominators are a fraction of example if you have (2/3)x + (3/4) = 6 the denominators are 3 and 4 (and the denominator of the 6 is 1, but you dont need to worry about it) the LCD of 3 and 4 is 12 12 is the smallest number divisible by BOTH 3 and 4 multiply everything by 12 to get rid of denominators

OpenStudy (laronjaxon):

omg thank you!!!!!!! i think i understand!!